OpenStudy (anonymous):

I am factoring trinomials and I need someone to refresh my memory on how to do it !! (: thanks

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (amistre64):

just try to remember how trinomial are formed; and usually it means you are dealing with a quadratic

OpenStudy (amistre64):

(x+b) (x+c) x+b x+c -------- x^2 +bx +cx +bc ---------------- x^2 +(b+c)x +(bc)

OpenStudy (amistre64):

to factor is to undo what you multiplied together

OpenStudy (amistre64):

x^2 +3x -10 for example; we need to find 2 numbers such that: (bc) = -10 and: (b+c) = 3

OpenStudy (amistre64):

(x+b) (x+c) is the factored form then: 1,10; -1(10) = -10, but -10 +1 not= 3 ; those aint workin' -2,5; -2(5) = -10; 5 -2 = 3; thatll work (x+5)(x-2)

OpenStudy (anonymous):

can you guys help me with\[a squared +4ab-45bsqaured\]